OpenStudy (anonymous):

Suppose y varies directly with x, and y=8 when x=-6. What direct variation equation relates x and y? What is the value of y when x=-2? A.y=-0.75x;1.50 B.y=-1.33x;2.67 C.y=1.33x;-2.67 D.y=0.13x;-0.25

OpenStudy (campbell_st):

\[y = -\frac{4}{3} x\] check by substituting \[y = -\frac{4}{3} \times (-6) = \frac{24}{3} = 8\] so true

OpenStudy (anonymous):

A

OpenStudy (campbell_st):

next , find y when x = 2 \[y = -\frac{4}{3} \times 2 = -8/3 = -2.67\]